OpenStudy (anonymous):

inverse of x^3+3x+1?

OpenStudy (anonymous):

solve x=y^3+3y+1

OpenStudy (anonymous):

oh also its inverse (1)...so plug in 1 im guessing?

OpenStudy (anonymous):

f^-1(1)

OpenStudy (anonymous):

yea but how do you solve it with two y's?

OpenStudy (anonymous):

if it is \[f^{-1}(1)\] you are solving \[x^3+3x+1=1\]

OpenStudy (anonymous):

it is very tough to solve it but if X matrix then u can solve it

OpenStudy (anonymous):

do i do inverse though? i dont understand

OpenStudy (anonymous):

no matrix hold on

OpenStudy (anonymous):

if \[f(a)=b\] then \[f^{-1}(b)=a\]

OpenStudy (anonymous):

so asking for \[f^{-1}(1)\] is the same as saying "what x will give an output of 1?"

OpenStudy (anonymous):

in other words, you know \[f(x)=1\] you just don't know x yet

OpenStudy (anonymous):

okay so...just tell me the first step i need to do? still a little confused

OpenStudy (anonymous):

wouldnt x=0?

OpenStudy (anonymous):

the first real step is to understand what \[f^{-1}(1)\] means. it is the number you plug in to \[f(x)\] to get 1 as an output. so your first algebra step is to write \[f(x)=1\] i.e. \[x^3+3x+1=1\] and solve for x

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

\[x^3+3x=0\] \[x(x^2+3)=0\] \[x=0\] or just do it in your head

OpenStudy (anonymous):

ohhh okay i see it

OpenStudy (anonymous):

i have another inverse question?

OpenStudy (anonymous):

great. solving \[f^{-1}(x)\] instead of \[f^{-1}(1)\] would be more algebra

OpenStudy (anonymous):

sure ask away

OpenStudy (anonymous):

okay f(x)=7x/(x+1). i found the inverse to be -x/(x-7). then it says to check using function composition? how do i do (f^-1 dot f)(x) and (f dot f^-1)(x)?

OpenStudy (anonymous):

a bunch of tedious algebra i will write the first one with explanation

OpenStudy (anonymous):

but before i do let me check your answer because i don't want to start and end up being wrong, hold on a second

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes good work! i might be easier if you wrote it as \[f^{-1}(x)=\frac{x}{7-x}\] but maybe not. we can work with yours

OpenStudy (anonymous):

now your job is to show that \[(f^{-1}\circ f)(x)=x\]

OpenStudy (anonymous):

first get rid of the circle and write what it means \[(f^{-1}\circ f)(x)=f^{-1}(f(x))\]

OpenStudy (anonymous):

then work from the inside out. replace the general f by the specific one you have and write \[f^{-1}(\frac{7x}{x+1})\]

OpenStudy (anonymous):

now compute by replacing the x in "f(x)" by 7x/(x+1) \[f^{-1}(\frac{7x}{x+1})=\frac{-(\frac{7x}{x+1})}{\frac{7x}{x+1}-7}\]

OpenStudy (anonymous):

yea i got this far but dont know what to do with the fractions in a fraction

OpenStudy (anonymous):

looks nice and ugly, but since the answer will just be x you should expect a whole raft of cancelation

OpenStudy (anonymous):

you can do it in three steps or less. first multiply numerator and denominator by x +1

OpenStudy (anonymous):

you get \[\frac{-7x}{7x-7(x+1)}\]

OpenStudy (anonymous):

now you are really done. just get \[\frac{-7x}{7x-7x-7}=\frac{-7x}{-7}=x\]

OpenStudy (anonymous):

i couldnt cancel out 7x/x+1 before?

OpenStudy (anonymous):

hell no!

OpenStudy (anonymous):

like saying \[\frac{5}{5+2}=\frac{\cancel{5}}{\cancel{5}+2}=\frac{1}{2}\]

OpenStudy (anonymous):

haha okay. so the next equation equals x too i see. what is the range of f is my last question from this problem?

OpenStudy (anonymous):

\[\frac{-(\frac{7x}{x+1})}{\frac{7x}{x+1}-7}\] cannot cancel anything until you write as one fraction

OpenStudy (anonymous):

range of f is the domain of f inverse lets me scroll up and see what f is

OpenStudy (anonymous):

\[f(x)=\frac{7x}{x+1}\] this ratio can never be exactly 7

OpenStudy (anonymous):

because of that 1 in the denominator

OpenStudy (anonymous):

and you can also see that the domain of f inverse is all real numbers except 7, because you cannot divide by zero

OpenStudy (anonymous):

ohhh okay i see! thank you i get it now

OpenStudy (anonymous):

in other words since no input for f will give an output of 7, 7 is not in the domain of f inverse. good luck!